OpenStudy (anonymous):

radical 49 divided by radical 2

OpenStudy (anonymous):

√49/√2=7/√2 √49=7

OpenStudy (karatechopper):

CONJUGANT! multiply by the conjugant! lets write this out numerically!

OpenStudy (anonymous):

this is what it looks like \[\sqrt{49} \div \sqrt{2}\]

OpenStudy (karatechopper):